# Time Series Information Visualization – A Review of Approaches and Tools

### v1

@@ 1 Introduction
The remainder of the paper is organized as follows: Section 2 presents the research methodology adopted for collecting literature; Section 3 presents previous reviews about InfoVis and the visualization of time series; Section 4 addresses the ground theory for the understanding of the time series domain and its applications; Section 5 is devoted to describing methods and approaches designed for time series InfoVis; Section 6 extends the previous section and presents composed interfaces applied to multidimensional time series exploration relying on multiple visualization approaches; Section 7 discusses open challenges and suggests future research directions; finally, Section 8 provides our final remarks.
@@ 2.1 Literature Search Procedure
[…]
@@ 3 Previous Visualization Reviews
Andrienko et al. [17] provided an early but comprehensive taxonomy of spatiotemporal visualization techniques, categorizing existing methods by data type, visual metaphor, and analytical tasks. However, the authors underestimated real-time interaction challenges. Recent works reviewed the complexities of visualizing spatiotemporal data [21, 178, 14]. Bai et al. [24] presented a systematic review of time-varying volumetric data frequently used in scientific visualization to understand the temporal evolution of complex phenomena such as fluid dynamics, cosmology, and climatology.
Despite the rich literature, effectively visualizing multidimensional time series remains a significant challenge. This work focuses specifically on solutions for time series information visualization (InfoVis). We provide an updated discussion on techniques and tools for interactive visual analysis that were previously unavailable or only recently emerged. Therefore, this study provides developers with a comprehensive overview, supporting the design of tools that enable data scientists to understand time-series data better and visually transform information into knowledge. In addition, we have made all the code and data used to render our visualizations publicly available in an online repository [^1] Available on GitHub at https://github.com/evortigosa/InfoVis.
@@ 4.1 Time Series Notation and Nomenclature
$D=\big{\{}(t_{1},\mathbf{v}_{1}),(t_{2},\mathbf{v}_{2}),\dotsc,(t_{n},\mathbf{v}_{n})\big{\}}$
being $t_{i}$ the instant of time and $\mathbf{v}_{i}=f\big{\langle}t_{i}\big{\rangle}$ the time-dependent vector (also called an instance or sample) that can bear properties of different types (e.g., an integer, a real, a categorical, an image, or a binary sequence) [189]. The term “time series” is frequently used to refer to time-oriented data comprising numerical values, while time-oriented data consisting of categorical values are sometimes termed event sequences [28]. To ensure terminological clarity and prevent misunderstandings, we refer to time-dependent data of any type as simple “time series”.
A time series dataset is considered unidimensional when it contains only a single qualitative or quantitative feature, also known as an attribute or variable. Although the temporal evolution of ordered pairs $(t_{i},\mathbf{v}_{i})$ can be represented in a Cartesian coordinate space, each instance of $\mathbf{v}_{i}$ can also bear a subset carrying an arbitrary number of features, expressed as $\mathbf{v}_{i}=\{v_{i1},v_{i2},\dotsc,v_{ik}\}$. These $k$ elements define the feature space of the time-dependent instances, thus leading to a multidimensional time series dataset, since each instance can be interpreted as a point in the $m$-dimensional space where $m=k+1$ is the total number of elements in $(t_{i},\mathbf{v}_{i})$, that is, the $k$ time-dependent features and the time component. Consequently, families of time-varying curves for each feature of time-dependent instances can be produced in the $m$-dimensional space [109].
@@ 4.2 Multidimensionality
Figure 1 illustrates the structure of a multidimensional time series dataset conceptually represented as a data cube. Consider, for example, a simulation procedure that involves parametric optimization, where a set of features is given as input for multiple simulation methods at each time step. More specifically, different vectors of input features are received and processed by different algorithms, and then multiple series of output vectors are produced. By recording both inputs and outputs at each time step, we obtain a snapshot of the simulation. Iterating this procedure over subsequent time instants for modeling features and possible adjustments to the simulation parameters generates a sequence of such snapshots that contains all stimuli and outcomes processed over time. When organizing these data as rows (series of outputs), columns (features), and depth (temporal evolution), we have a multidimensional time series that resembles a cube-like temporal structure, as shown in Figure 1.
As the dimensionality of the “time cube” increases, representing and interpreting its content becomes increasingly challenging, as it may present time-varying groups of relational structures. This increased complexity complicates the analytical tasks related to extracting useful information from the data. In particular, analyzing trends and interdependencies among time-varying features is at the core of multidimensional time series mining [77].
@@ 4.3 Overlook Through Time Series Modeling
where the combination of cosine and sine functions is alongside functions $A(\cdot)$ and $B(\cdot)$ in terms of differentials indexed on the continuous parameter $\omega$, easily extended from $[0,\pi]$ to $[-\pi,\pi]$ domain, allowing the existence of the spectral distribution/density function. There are standard methods of frequency transformation (FT), such as the discrete Fourier transform (DFT), discrete cosine transform (DCT), or discrete wavelet transform (DWT), which can be used to detect periodicities and frequency content. Recent studies integrate spectral coefficients and ANN models to uncover deeper frequency-based patterns [73, 198].
In practice, time-domain components can be visualized through coordinated views, supporting the dynamic exploration of each element. Frequency domain visualizations often focus on spectral plots of dominant periods, where the number of observations between seasonal repetitions (the “frequency”) guides interpretation.
@@ 4.4 Time Series and Information Visualization
The final stage involves user interaction. Using an input device, such as a computer mouse, analysts select intervals of interest to explore through comparisons among different data views, filtering out noise, and converting the visualized information into actionable knowledge. Although information tends to be cumulative, knowledge is selective. Only the information displayed on a computer screen may not be sufficient for making a decision, and the excess of information can clutter the analysis. Therefore, when performing an analysis, the analysts must know what to do with the discovered information, since it is the foundation of knowledge.
@@ 4.6 Graphical Perception and Visual Properties
Key perceptual properties that can guide the graphical encoding include: ($i$) groups of units sharing similar goals (e.g., common fate or proximity); ($ii$) groups of elements with similar shape characteristics (e.g., size, color, orientation, symmetry, or parallelism); ($iii$) groups of objects related by the same action; ($iv$) smooth or curving elements perceived as more visually friendly; and ($v$) groups of elements enclosed within well-delimited boundaries. Figure 5 illustrates aesthetic principles based on Gestalt theory.
@@ 4.7 Interactive Exploratory Techniques
Multi-view systems Simultaneously display multiple coordinated visualizations to support comparison, cross-referencing, and exploration across different aspects of time-series data without forcing context switches.
Overview first, zoom and filter, then details-on-demand A visual exploration paradigm where users start with a high-level temporal summary to identify macro-patterns, refine intervals of interest through zooming and filtering, and finally access detailed data or segments selectively.
@@
Zoom Interactive scaling of a selected time interval. Expands the temporal axis to reveal intra-interval detail, allowing the inspection of fine-grained variations in the data.
@@
Tooltips On-hover text pop-ups displaying timestamped values, metadata, or analytic annotations. Enable value inspection without requiring persistent visual changes.
@@
5 Approaches and Methods
@@ 6.1 Textual Documents and Topics
Specifically, the interface includes four main components: the EvoRiver chart itself, describing the flow of topics’ evolution over time; a bivariate trend chart showing variations in topics’ behaviors (positive/negative competition/collaboration power); edges linking related topics; and a Word Cloud [88] presenting keywords for a selected topic, with word size proportional to frequency. Sun et al. [166] demonstrated the interface using Twitter data to analyze the evolution of topics discussed during the 2012 US presidential election. They found that the topics “Government” and “Politics” were often closely related, but under the influence of “International issues”, “Government” and “Politics” alternated between competitors and collaborators, thus allowing analysts to understand the complex dynamics of the main topics covered during that election process [42]. EvoRiver relies on a sentiment classifier based on statistical methods and expert labeling, which may introduce subjective bias to the technique.
@@ 6.2 Graph-structured Time Series
Visualizing graph-structured time series for large datasets remains challenging due to the computational demands of rendering their node-link structures and the tendency for occlusions. Therefore, choosing an efficient layout algorithm is critical to mitigate overlaps (force-directed with edge bundling or hierarchical edge routing). Interactive tools such as zooming or fisheye distortion could be used to navigate between connections, in addition to color encodings to highlight clusters (see Section 4.7).
@@
Later, Nascimento et al. [134] summarized more than 1.5 billion 256-channel electroencephalogram (i.e., 256 time series) observations into 15 million symbolic intervals (approximately 0.98% of the original size) through dynamic linear modeling, discovering temporal patterns beyond simple average trends. Moreover, questions of dynamic accommodation, as observed in time series intervals in neuroscience research, were revealed through visualizations of brain activity patterns using an interactive graph that highlighted the group communication of the ROIs across time, also obtained through quantile regression on symbolic intervals.
@@ 6.4 Environmental Monitoring
The authors designed a bivariate chart between the top and bottom groups to link their information, and provide an intuition of the deviation between the two datasets. It is based on the statistical principle of bivariate analysis and visually contrasts two time series using an interpolation-based strategy that creates two concurrent curves that highlight the differences with contrasting colors. The tool is a coordinated multi-view interface that supports interactive brushing and zooming, and encodes multiple information channels simultaneously: seasonal color coding, tooltips, PM10 concentration averages, tooltips, and dynamically updated legends.
Multidimensional (or even high-dimensional) time series are ubiquitous in Industry 4.0, and alternative Symbolic Data Analysis (SDA) representations for time series have emerged. Fonseca et al. [64] applied a symbolic representation to weather data from water-particle monitoring in Chile’s Atacama Desert. Using a statistical process control approach, they proposed a bivariate SDA control chart as a daily humidity monitoring tool. Given their flexibility and speed in data convergence, SDA-based models can incorporate dynamic events related to complex time-dependent data, standing as computationally efficient alternatives for summarizing, detecting outliers, and visualizing high-dimensional time series.
@@ 7.1 Software Tools and Libraries for InfoVis
The creation of InfoVis solutions has become increasingly accessible due to the availability of high-quality open-source libraries in multiple programming languages. Matplotlib (https://matplotlib.org/) is a Python library that provides a wide range of highly customizable plotting functions, supported by extensive documentation and numerous implementation examples suitable for both beginners and experienced developers. Similarly, Seaborn (https://seaborn.pydata.org/) and Plotly (https://plotly.com/python/) are valuable Python libraries used to generate complex, publication-quality visualizations. Datashader (https://datashader.org/) is an open-source Python library specifically designed to visualize very large datasets.
$\mathbf{R}$ is a statistical programming language well-known for its powerful data processing capabilities. For producing visualizations in $\mathbf{R}$, we refer to the ggplot2 grammar-based package [190]. JavaScript is the standard programming language for web applications and enables the creation of interactive and responsive HTML5 objects [78]. In this context, D3.js [31] is a well-documented JavaScript library that enables dynamic data mapping and interactive visualizations using canvas and SVG graphical objects [132]. D3.js allows visualization developers to create highly customizable and interactive interfaces; however, it demands a solid experience in web development with JavaScript. NetworkD3 (https://christophergandrud.github.io/networkD3/) is an $\mathbf{R}$ package that leverages D3.js to create interactive graphs rendered with HTML compatibility, bridging $\mathbf{R}$ programming with web applications.
Tableau (https://www.tableau.com/) and Microsoft Power BI (https://powerbi.microsoft.com) are advanced business intelligence software tools used for data analysis and the creation of interactive dashboards through a variety of interactive visualization techniques. However, the full-featured versions of these tools can be expensive.
@@
Several other solutions, ranging from advanced platforms to easy-to-use and straightforward options, can be found online for chart generation. Here, we focus on the most popular and well-documented tools and libraries.

### v2

@@ 1 Introduction
1.1 Organization
The remainder of the paper is organized as follows: Section 2 presents the research methodology adopted for collecting literature; Section 3 presents previous reviews covering different aspects of visualization for time series; Section 4 addresses the ground theory to understand the time series domain and its applications, which includes formal definitions (4.1), dimensionality (4.2), missing data (4.3), modeling approaches (4.4), an introduction to the visualization pipeline (4.5), taxonomies and guidelines (4.6), graphical properties (4.7), interactivity (4.8), and large-scale datasets (4.9); Section 5 describes visualization methods and approaches designed for time series InfoVis; Section 6 extends the previous section and presents composed interfaces applied to multidimensional time series in the context of text and topics (6.1), graphs (6.2), financial data (6.3), environmental monitoring (6.4), audio and images (6.5), and event sequences (6.6); Section 7 discusses software tools for time series InfoVis (7.1), open challenges (7.2), and suggests future research directions (7.3); finally, Section 8 provides our final remarks.
@@ 2.1 Literature Search Procedure
Papers that did not satisfy at least one of the selection criteria were excluded. The queries returned 297 publications, and duplicates were removed after a second filtering based on abstract and introduction readings. For each selected paper, we reviewed the main content and did not consider the appendices or supplementary material. A total of 232 studies were then read in full and, after a careful review of their contents, ???? references are reported in this review.
@@ 3 Previous Visualization Reviews
Andrienko et al. [17] provided an early but comprehensive taxonomy of spatiotemporal visualization techniques, categorizing existing methods by data type, visual metaphor, and analytical tasks. However, the authors underestimated real‑time interaction challenges. Recent works reviewed the complexities of visualizing spatiotemporal data [21, 191, 14]. Bai et al. [24] presented a systematic review of time-varying volumetric data frequently used in scientific visualization to understand the temporal evolution of complex phenomena such as fluid dynamics, cosmology, and climatology.
Despite the rich literature, effectively visualizing multidimensional time series remains a significant challenge. This work focuses specifically on solutions for time series information visualization (InfoVis). We provide an updated discussion on techniques and tools for interactive visual analysis that were previously unavailable or only recently emerged. Therefore, this study provides developers with a comprehensive overview, supporting the design of tools that enable data scientists to better understand time-series data and visually transform information into knowledge. In addition, we have made all the code and data used to render our visualizations publicly available in an online repository [^1] Available on GitHub at https://github.com/evortigosa/InfoVis.
@@ 4.1 Time Series Notation and Nomenclature
$D=\big\{(\mathbf{t}_{1},\mathbf{v}_{1}),(\mathbf{t}_{2},\mathbf{v}_{2}),\dotsc,(\mathbf{t}_{n},\mathbf{v}_{n})\big\}$
being $t_{i}$ the instant of time and $\mathbf{v}_{i}=f\big\langle\mathbf{t}_{i}\big\rangle$ the time-dependent vector (also called an instance or sample) that can bear properties of different types (e.g., an integer, a real, a categorical, an image, or a binary sequence) [202]. The term “time series” is frequently used to refer to time-oriented data comprising numerical values, while time-oriented data consisting of categorical values is sometimes termed event sequences [28]. To ensure terminological clarity and prevent misunderstandings, we refer to time-dependent data of any type as simple “time series”.
A time series dataset is considered unidimensional when it contains only a single qualitative or quantitative feature, also known as an attribute or variable. Although the temporal evolution of ordered pairs $(\mathbf{t}_{i},\mathbf{v}_{i})$ can be represented in a Cartesian coordinate space, each instance of $\mathbf{v}_{i}$ can also bear a subset carrying an arbitrary number of features, expressed as $\mathbf{v}_{i}=\{v_{i1},v_{i2},\dotsc,v_{ik}\}$. These $k$ elements define the feature space of the time-dependent instances, thus leading to a multidimensional time series dataset, since each instance can be interpreted as a point in the $m$-dimensional space where $m=k+1$ is the total number of elements in $(\mathbf{t}_{i},\mathbf{v}_{i})$, that is, the $k$ time-dependent features and the time component. Consequently, families of time-varying curves for each feature of time-dependent instances can be produced in the $m$-dimensional space [114].
@@ 4.2 Multidimensionality
Figure 1 illustrates the structure of a multidimensional time series dataset conceptually represented as a data cube. Consider, for example, a simulation procedure that involves parametric optimization, where a set of features is given as input for multiple simulation methods at each time step. More specifically, different vectors of input features are received and processed by different algorithms, and then multiple series of output vectors are produced. By recording both inputs and outputs at each time step, we obtain a snapshot of the simulation. Iterating this procedure over subsequent time instants for modeling features and possible adjustments to the simulation parameters generates a sequence of such snapshots that contains all stimuli and outcomes processed over time. When organizing these data as rows (series of outputs), columns (features), and depth (temporal evolution), we have a multidimensional time series that resembles a cube-like temporal structure, as shown in Figure 1, which can be extended by adding spatial features or stacking temporal layers into the data structure.
As the dimensionality of the “time cube” increases, representing and interpreting its content becomes increasingly challenging [144], as it may present time-varying groups of relational structures. This increased complexity complicates the analytical tasks related to extracting useful information from the data. In particular, analyzing trends and interdependencies among time-varying features is at the core of multidimensional time series mining [81].
4.3 Fragmented and Intermittent Time Series
Real-world time series often contain noise and interruptions due to sensor failure, communication loss, or irregular sampling intervals. These interruptions produce missing observations with distinct statistical characteristics that can obscure the underlying patterns [59], which require explicit treatment to avoid biased inference while preserving the integrity of the temporal dynamic.
Approaches for handling fragmented time series fall into two broad categories. First, classical statistical methods, such as linear or spline interpolation, ARIMA and Kalman filters, and state-space methods, estimate missing data points when the missing spans are short and the underlying dynamics are locally smooth [120, 156]. For example, Alexandrina et al. [10] employed a spline interpolation approach to represent fragmented points in an interactive system that dynamically adjusts the visualizations. Second, modern data-driven machine learning techniques directly learn temporal and cross-series dependencies and can impute long or structured gaps [192]. For example, Cao et al. [37] proposed a bidirectional recurrent model that propagates information in forward and backward directions, demonstrating effectiveness in reconstructing missing extended segments.
From the visualization perspective, designers must balance continuity and transparency, i.e., communicating data imputation. Best practices can include: exposing the missingness pattern and provenance (so analysts know which values were measured versus inferred) [175]; rendering imputed values together with explicit uncertainty encodings, such as error bars or translucent bands [96]; and providing interactive controls that allow users to switch between raw and imputed views, to view per-segment confidence intervals, and to inspect the imputation method and parameters on demand.
@@ 4.4 Overlook Through Time Series Modeling
where the combination of cosine and sine functions is alongside functions $A(\cdot)$ and $B(\cdot)$ in terms of differentials indexed on the continuous parameter $\omega$, easily extended from $0,\pi]$ to $\pi,\pi]$ domain, allowing the existence of the spectral distribution/density function. There are standard methods of frequency transformation (FT), such as the discrete Fourier transform (DFT) or discrete cosine transform (DCT), which can be used to detect periodicities and frequency content.
Real-world time series often arise as wave-like patterns in natural and engineered experiments, embedding concepts from physics such as energy transfer, phase shifts, and oscillatory dynamics. In this context, wavelet transforms decompose non-stationary series into time-frequency components, localizing energy transfer and phase changes that Fourier transforms can overlook, enabling time series visualization as oscillatory waves. Cazelles et al. [40] applied the continuous wavelet transform (CWT) to generate decompositions that identify transient, non-stationary oscillatory behavior in epidemiological time series. Priyadarshini et al. [157] used wavelet packet decomposition (WPD) to detect and visualize voltage disturbances in a computationally efficient power quality monitoring approach.
In practice, time-domain components can be visualized through coordinated views, supporting the dynamic exploration of each element. Frequency domain visualizations often focus on spectral plots of dominant periods, where the number of observations between seasonal repetitions (the “frequency”) guides interpretation. Recent studies have integrated spectral coefficients and ANN models to uncover deeper frequency-based patterns [77, 211].
@@ 4.5 Time Series and Information Visualization
The final stage involves user interaction. Using an input device, such as a computer mouse, analysts select intervals of interest to explore through comparisons among different data views, filtering out noise, and converting the visualized information into actionable knowledge. Although information tends to be cumulative, knowledge is selective. Only the information displayed on a computer screen may not be sufficient for making a decision, and an excess of information can clutter the analysis. Therefore, when performing an analysis, the analysts must know what to do with the discovered information, since it is the foundation of knowledge.
@@ 4.7 Graphical Perception and Visual Properties
Key perceptual properties that can guide the graphical encoding include: ($i$) groups of units sharing similar goals (e.g., common fate or proximity); ($ii$) groups of elements with similar shape characteristics (e.g., size, color, orientation, symmetry, or parallelism); (iiiiii) groups of objects related by the same action; ($iv$) smooth or curving elements perceived as more visually friendly; and ($v$) groups of elements enclosed within well-delimited boundaries. Figure 5 illustrates aesthetic principles based on Gestalt theory.
@@ 4.8 Interactive Exploratory Techniques
Multi-view systems Simultaneously display multiple coordinated visualizations to support comparison, cross-referencing, and exploration across different aspects of time‑series data without forcing context switches.
Overview first, zoom and filter, then details-on-demand A visual exploration paradigm where users start with a high‑level temporal summary to identify macro‑patterns, refine intervals of interest through zooming and filtering, and finally access detailed data or segments selectively.
@@
Zoom Interactive scaling of a selected time interval. Expands the temporal axis to reveal intra‑interval detail, allowing the inspection of fine-grained variations in the data.
@@
Tooltips On‑hover text pop‑ups displaying timestamped values, metadata, or analytic annotations. Enable value inspection without requiring persistent visual changes.
@@
4.9 Large-Scale and Real-Time Time Series
The proliferation of big data has generated extreme-scale time series containing millions or even billions of samples, requiring techniques capable of handling such a large volume of data while keeping computational efficiency, perceptual clarity, and interactivity. Two complementary classes of methods are commonly used in this context: aggregation/downsampling that reduces raw samples to a compact representation and multi-view systems that progressively refine visual detail on demand (see Section 4.8).
A straightforward aggregation is temporal binning, where data points are grouped into fixed intervals (e.g., hourly or daily bins) and summarized using averages, medians, or densities, which reduces data volume while preserving high-level trends [42]. For example, imMens [130] employed binned aggregation for the real-time querying of massive datasets, allowing subsecond response times for interactive explorations, while TiVy [42] provides visual summaries through adaptive binning and summarization. Similarly, curve-preserving downsampling algorithms select representative points to conserve salient visual features. In this context, M4 [106] performs query-level aggregation to produce error-bounded line visualizations at very high reduction rates.
Multidimensional (or even high-dimensional) time series are ubiquitous in Industry 4.0, and alternative Symbolic Data Analysis (SDA) representations for time series have emerged. Nascimento et al. [142] summarized more than 1.5 billion 256-channel electroencephalogram (i.e., 256 time series) observations into 15 million symbolic intervals (approximately 0.98% of the original size) through dynamic linear modeling, discovering temporal patterns beyond simple average trends. Moreover, questions of dynamic accommodation, as observed in time series intervals in neuroscience research, were revealed through visualizations of brain activity patterns using an interactive graph that highlighted the group communication of the brain’s Regions of Interest (ROIs) across time, also obtained through quantile regression on symbolic intervals. Given their flexibility and speed in data convergence, SDA-based models can incorporate dynamic events related to complex time-dependent data, demonstrating computational efficiency to summarize, detect outliers, and visualize high-dimensional time series.
In real-time contexts, systems must process data in (near) constant time. Progressive partitioning and multidimensional pattern extraction incrementally update aggregations through operations that bin data on-the-fly [126], while deep learning-based analytics approaches integrate visual analytics with neural networks to handle massive datasets by extracting embeddings for clustered visualization [162]. Refer to Ulmer et al. [186] for a comprehensive overview of progressive visualization and visual analytics.
5 Visual Approaches and Methods
@@ 6.1 Textual Documents and Topics
Specifically, the interface includes four main components: the EvoRiver chart itself, describing the flow of topics’ evolution over time; a bivariate trend chart showing variations in topics’ behaviors (positive/negative competition/collaboration power); edges linking related topics; and a Word Cloud [92] presenting keywords for a selected topic, with word size proportional to frequency. Sun et al. [178] demonstrated the interface using Twitter data to analyze the evolution of topics discussed during the 2012 US presidential election. They found that the topics “Government” and “Politics” were often closely related, but under the influence of “International issues,” “Government,” and “Politics” alternated between competitors and collaborators, thus allowing analysts to understand the complex dynamics of the main topics covered during that election process [45]. EvoRiver relies on a sentiment classifier based on statistical methods and expert labeling, which may introduce subjective bias to the technique.
@@ 6.2 Graph-structured Time Series
Visualizing graph-structured time series for large datasets remains challenging due to the computational demands of rendering their node-link structures and the tendency for occlusions. Therefore, choosing an efficient layout algorithm is critical to mitigate overlaps (force‑directed with edge bundling or hierarchical edge routing). Interactive tools such as zooming or fisheye distortion could be used to navigate between connections, in addition to color encodings to highlight clusters (see Section 4.8).
@@ 6.4 Environmental Monitoring
The authors designed a bivariate chart between the top and bottom groups to link their information and provide an intuition of the deviation between the two datasets. It is based on the statistical principle of bivariate analysis and visually contrasts two time series using an interpolation-based strategy that creates two concurrent curves that highlight the differences with contrasting colors. The tool is a coordinated multi-view interface that supports interactive brushing and zooming, and encodes multiple information channels simultaneously: seasonal color coding, tooltips, PM10 concentration averages, and dynamically updated legends.
Similarly, Fonseca et al. [68] applied a symbolic representation to weather data from water-particle monitoring in Chile’s Atacama Desert. Using a statistical process control approach, they proposed a bivariate SDA (see Section 4.9) control chart as a daily humidity monitoring tool.
@@ 7.1 Software Tools and Libraries for InfoVis
The creation of InfoVis solutions has become increasingly accessible due to the availability of high-quality open-source libraries in multiple programming languages. Matplotlib (https://matplotlib.org/) is a Python library that provides a wide range of highly customizable plotting functions, supported by extensive documentation and numerous implementation examples suitable for both beginners and experienced developers. Similarly, Seaborn (https://seaborn.pydata.org/) and Plotly (https://plotly.com/python/) are valuable Python libraries used to generate complex, publication-quality visualizations. Datashader (https://datashader.org/) is an open‐source Python library specifically designed to visualize very large datasets.
$\mathbf{R}$ is a statistical programming language well-known for its powerful data processing capabilities. For producing visualizations in $\mathbf{R}$, we refer to the ggplot2 grammar-based package [203]. JavaScript is the standard programming language for web applications and enables the creation of interactive and responsive HTML5 objects [82]. In this context, D3.js [31] is a well-documented JavaScript library that enables dynamic data mapping and interactive visualizations using canvas and SVG graphical objects [140]. D3.js allows visualization developers to create highly customizable and interactive interfaces; however, it demands a solid experience in web development with JavaScript. NetworkD3 (https://christophergandrud.github.io/networkD3/) is an $\mathbf{R}$ package that leverages D3.js to create interactive graphs rendered with HTML compatibility, bridging $\mathbf{R}$ programming with web applications. Shiny (https://shiny.posit.co/) is an open source package that integrates with ggplot2 from $\mathbf{R}$ and Seaborn/Plotly from Python, allowing users to create high-quality web applications for time series visualization based on a reactive programming model, which supports real-time updates and dynamic user interactions.
Tableau (https://www.tableau.com/) and Microsoft Power BI (https://powerbi.microsoft.com) are advanced business intelligence (BI) software tools used for data analysis and the creation of interactive dashboards using a variety of interactive visualization techniques. However, full-featured versions of these tools require proprietary licensing and can be expensive.
@@
Table 4 summarizes the strengths and limitations of each tool described in this section. Several other solutions, ranging from advanced platforms to straightforward and easy-to-use options, can be found online for chart generation. Here, we focus on widely used and well-documented tools and libraries.
Table 4: Summary of Tools and Libraries for InfoVis.
Tool Description Strengths Limitations Interactivity
Matplotlib Python library for customizable plotting functions. Highly customizable; suitable for publication-quality figures. Complex coding for advanced features; not optimized for very large data. Low (static by default).
Seaborn Python library built on Matplotlib for statistical visualizations. Simplifies creation of complex plots compared to Matplotlib. Similar performance limitations of Matplotlib with large data. Low to moderate (static focus).
Plotly Python library for interactive, web-ready visualizations. Enables dynamic and shareable visuals; web integration. Complex coding for advanced features. High (provides zoom and click events).
Datashader Python library for visualizing very large data. Efficient for massive volumes; produces high-fidelity images. Less intuitive for small data. Moderate (can integrate with interactive tools).
ggplot2 Grammar-based $\mathbf{R}$ package. High-capable for statistical graphics; extensible. $\mathbf{R}$-specific. Low (static by default).
D3.js JavaScript library for dynamic data mapping. SVG-native; flexible for custom interfaces; supports animations. Requires strong skills in web development. Very high (focus on interactivity).
NetworkD3 $\mathbf{R}$ package leveraging D3.js for interactive graphs. Bridges $\mathbf{R}$ coding with web interactivity. Small number of visualizations; requires $\mathbf{R}$ and JavaScript skills. High (D3.js-based).
Shiny Web app for data analysis and interactive dashboards. User-friendly; integrates with $\mathbf{R}$ or Python ecosystems. Complex coding for advanced features; free version is limited. High (supports brushing, zooming, sliders).
Tableau BI software for data analysis and interactive dashboards. User-friendly; dashboards with real-time updates. Full features are expensive; proprietary licensing. Very high (built-in dashboards).
Power BI BI software for data analysis and interactive dashboards. User-friendly; supports collaboration. Full features are expensive; tied to Microsoft environment. Very high (built-in dashboards).
General-Purpose LLMs Generative models for textual content and code. Accelerates prototyping; helps with error checking in coding. Outputs require strong verification; potential inaccuracies. Text-based interaction for queries and answers.
Specialized LLMs Generative models focused on visualization code synthesis. Automates development from natural language. Lack maturity (emerging field). High (interactive code generation).